OpenStudy (toxicsugar22):

(x-4)^2/9-(y+2)^2/16=1 give the exact coordinates of any four points on the hyperbola. and show me how you get this

OpenStudy (anonymous):

wouldn't you just plug in any x or y number and factor to get the corresponding x or y value?

OpenStudy (anonymous):

First, try setting x=0, then y=0, these should be the easiest 2 points to get

OpenStudy (anonymous):

agreed

OpenStudy (toxicsugar22):

yes but can u help me further i tried

OpenStudy (toxicsugar22):

hi can u help me

OpenStudy (anonymous):

do you have a mac?

OpenStudy (anonymous):

if so, open grapher and plot the equation, this will let you see the curve.

OpenStudy (anonymous):

here is the algebra for x=0 \[16/9 - (y+2)^{2}/16 = 1\] \[5/9 = (y+2)^{2}/16\] \[40/9 = (y+2)^{2}\] \[y = -2+ (\sqrt{40})/3\]

OpenStudy (anonymous):

so that coordinate would be \[(0, -2+(\sqrt{40})/ 3)\]